Grade 10Electric Current

How to determine potential gradient of a given wire?

Arka Ghosh
9 Years agoGrade 10

1 Answer

Umakant biswal
9 Years ago
@arka 
potential gradient is the drop in potential per unit length of a wire . 
so, acc to the formula 
potential gradient = R/L 
so, as R= RHO*L/A 
SO, BY modifying potential gradient will be 
(rhol/a ) /L = rho / a 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..